please answer and show all steps

(1) Solve the right triangle where A = 40°, b= 6 (2) Solve the right triangle where a = 7, b= 11
Answer #1

(1) A=40°, 6=6. eight angle triangle - B - 90 C= 180 - (48 +90°) = c = 50° sinA Sin B Sinc b c - Sin 400 a Singoo 6 Sim 50° с